Question 21 (5 points) Tim Sport bought four bonds of ABC 6 1/2 16 at 78.125. Can you help Tim calculate his (A) total annual interest, (B) total cost, and (C) current yield to the nearest tenth percent?
Question 21 (5 points) Tim Sport bought four bonds of ABC 6 1/2 16 at 78.125. Can you help Tim calculate his (A) total annual interest, (B) total cost, and (C) current yield to the nearest tenth percent?
Answer #1

1.
Total annual interest=6.5%*100*4=26

2.
Total cost=78.125*4=312.50

3.
Current yield=6.5/78.125=8.320%